# PAID: A Framework of Product-Centric Advertising Image Design

### v1

@@ 4.4. Background Image generation
In this process, we train an SDXL-based (Podell et al., 2024) image generation model controlled by product foreground, prompt, and layout.
Layout control. Suppose there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. Supposing there are $M$ visual embeddings (V) after the self-attention layer. The output of the GSA layer is
$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$
, where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
Considering that layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. This reduces the parameters and inference latency, and has no negative impact on generation quality. Additionally, we adopt a LoRA Adaptation Training (LAT) strategy that trainable LoRA (Hu et al., 2022) layers are added to UNet, for bridging the gap between the base model and added GSA layers. This further accelerates convergence and increases generation quality. Figure 6 describes how layout control is achieved.
Figure 6. Illustration of layout control.
Training stages. To build the product-reserving layout-to-image model, the training involves two steps. First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it for deriving a layout-to-image model. We use a multi-scale training strategy for faster convergency, first training on 512x512 resolution images and then fine-tuning on 1024x1024 resolution images. Second, we equip the trained model with an inpainting ControlNet and further fine-tune it on 1024x1024 resolution images. This step incorporates the product foreground, achieving full control over the prompt, layout, and product foreground.
@@ 4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of text and other graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering. We first sort the text elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024). The selection is based on the similarity to the foreground color and contrast with the background area. Texts grouped by size and position use the same font and color. For the underlay, the color is chosen based on its contrast with the text color and similarity to the background color. The shape is selected from a predefined SVG library based on its size and proportions, with minor adjustments for fitting.
@@ 5.1. Implementation Details
For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone. Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images. The LLM is based on InternLM2 (Cai et al., 2024). We fine-tune the LLM branch on the PITA dataset for one epoch. We train the model using AdamW optimizer with a learning rate of 1e-5. Batch size on each GPU is set to 1 and gradient accumulation of 8 is adopted. The training costs about 14 hours with 8 NVIDIA A100 GPUs. During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use SDXL base model as the backbone. During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used. We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution. Then in step 2, we introduce an inpainting ControlNet which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
@@ 5.2. Evaluation Metrics
Overall pipeline. We evaluate the effectiveness of the overall pipeline from two perspectives: overall visual quality and layout. For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score. The aesthetic score is obtained by advertising experts ranking the outputs of different pipelines. The average rank is calculated to represent the mean aesthetic rating, where a lower rank indicates superior creative results. For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
Prompt generation. We evaluated prompt generation quality based on fore-background matching and e-commerce domain score.
  • FBM Rate (Fore-background matching rate): It refers to how well a product foreground fits with a given prompt background. We classify this matching into “reasonable” or “unreasonable” through human annotation, and then determine the proportion of reasonable cases for each method.
  • ED Score (E-commerce domain score): To assess how well the model-generated prompts reflect e-commerce characteristics, we calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Val means the ratio of elements larger than 0.1% of the canvas. Ali assesses how well elements are spatially aligned. Ove is the average intersection over union (IoU) of all element pairs except for underlays. $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements. Content metrics assess the harmony of layout with the background image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. Uti evaluates space usage for graphic elements. Occ is the average overlapping area between the graphic elements and products. Rea represents the non-flatness of regions that text elements without underlays are put on. For methods which take taglines as input, we report the text match rate (TMR), to see if the number of generated text bboxes are consistent with input taglines.
Background image generation. We evaluate the quality of generated background images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2023). To verify layout control, we use Grounding-DINO (Liu et al., 2023b) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth. The average of these maximum IoUs is called mIOU. In addition, Grounding-DINO AP scores are also reported.
@@ 5.3. Comparison
Compare with automatic pipelines for product advertising image design. Since there are no public automatic methods for advertising image design conditioned on the product foreground, taglines and target size, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation based on the input conditions. We fine-tune the SDXL-based inpainting model on our PIL dataset to obtain the image generation model. The only difference between this method and ours is the absence of layout control and the corresponding training stage. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt according to the foreground image using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general athetic rules (see details in the Appendix). Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison between these pipelines and our PAID framework. PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 7). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded text boxes, making them hard to read, while our model provides more flexible and readable layouts.
@@
Figure 7. Visualization of advertising image designed by different methods.
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score. We also use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 8). This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
Table 2. Evaluation of prompt generation methods
@@
Figure 8. T-sne visualization of generated prompts.
Compare with layout generation methods. To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a). For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased set and the generated set. The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased area. The generated set contain images re-generated using their ground-truth prompts (image captions). The generated set has no above hints and conforms to real applications. As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and product foreground. Our method uses the ground-truth prompt, product foreground, and taglines in PITA for prediction. The results are shown in Table 3 and Figure 9. Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements. This may be due to the fixed background image content, leaving less room for the model flexibility. RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating text boxes that cannot adequately fill the tagline content. The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images. Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
Table 3. Quantitative performance comparison of layout generation methods on PITA dataset. Bold number represents the best result in each column except for erased set. BG-based and FG-based are the abbreviations for Background-based and Foreground-based methods, respectively.
@@
Figure 9. Qualitative comparison of layout generation methods on PITA test set. Orange: text, Yellow: underlay, Red: logo, Black: Invalid
@@
5.4. Ablation Studies
Effect of JPGNL in layout generation. Our method predicts the overall layout of graphic and nongraphic elements. Here, we examine the need to predict the layout of nongraphic elements according to the background prompt. As shown in Figure 11, without specifying the position of nongraphic elements, texts placed on complex areas can lead to a messy appearance. By predicting the overall layout, we improve the text readability of advertising images. Additionally, with nongraphic layout predictions, generated images align more closely with background prompts. Without JPGNL, images might miss some objects from the prompts. Besides, in Table 5, we quantitatively analyze the impact of predicting nongraphic layouts on graphic layouts. Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and text readability. However, since it adds complexity to the task, it negatively affects graphic metrics such as element overlap and alignment.

### v2

@@ 4.4. Background Image generation
In this process, we train an SDXL-based (Podell et al., 2023) image generation model controlled by product foreground, prompt, and layout.
Layout control. Suppose that there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. Supposing there are $M$ visual embeddings (V) after the self-attention layer. The output of the GSA layer is
$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$,
where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
Considering that the layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. This reduces the parameters and inference latency, and has no negative impact on generation quality. Additionally, we adopt a LoRA Adaptation Training (LAT) strategy in which trainable LoRA (Hu et al., 2021) layers are added to UNet, to bridge the gap between the base model and added GSA layers. This further accelerates convergence and increases generation quality. Figure 4 describes how layout control is achieved.
Training stages. To build the product reservation layout-to-image model, training involves two steps. First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it to derive a layout-to-image model. We use a multi-scale training strategy for faster convergency, first training on 512x512 resolution images and then fine-tuning on 1024x1024 resolution images. Second, we equip the trained model with an inpainting ControlNet and further fine-tune it on 1024x1024 resolution images. This step incorporates the product foreground, achieving full control over the prompt, layout, and product foreground.
@@ 4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering. We first sort the tagline elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024b). The selection is based on the similarity to the foreground color and contrast with the background area. Taglines grouped by size and position use the same font and color. For the underlay, the color is chosen based on its contrast with the tagline color and similarity to the background color. The shape is selected from a predefined Scalable Vector Graphics (SVG) library based on its size and proportions, with minor adjustments for fitting.
@@ 5.1. Implementation Details
For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone. Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images. The LLM is based on InternLM2 (Cai et al., 2024). We fine-tune the LLM branch on the PITA dataset for one epoch. We train the model using AdamW optimizer with a learning rate of 1e-5. Batch size on each GPU is set to 1, and gradient accumulation of 8 is adopted. The training costs about 14 hours with 8 NVIDIA A100 GPUs. During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use the SDXL base model as the backbone. During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used. We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution. Then in step 2, we introduce an inpainting ControlNet, which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
@@ 5.2. Evaluation Metrics
To evaluate the overall poster generation and each module, we use the following metrics.
Overall pipeline. We evaluate the overall pipeline from three perspectives: overall visual quality, and layout. For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score. The aesthetic score comes from advertising experts ranking the outputs of different methods. The average rank represents the mean aesthetic rating, with a lower rank indicating better creative results. For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
Prompt generation. We evaluate prompt generation quality based on fore-background matching rate (FBM rate) and e-commerce domain score (ED score). FBM Rate refers to how well a product foreground fits with a given prompt background. We classify this matching into “reasonable” or “unreasonable” through human annotation, and calculate the proportion of reasonable cases. ED Score assesses how well the model-generated prompts reflect e-commerce characteristics. We calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
Figure 5. Visualization of advertising images designed by different methods.
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Val means the ratio of elements larger than 0.1% of the canvas. Ali assesses how well elements are spatially aligned. Ove is the average intersection over union (IoU) of all element pairs except for underlays. $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements. Content metrics assess the harmony of layout with the image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. Uti evaluates space usage for graphic elements. Occ is the average overlapping area between the graphic elements and products. Rea represents the non-flatness of regions that tagline elements without underlays are put on. For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes is consistent with input taglines.
Background image generation. We evaluate the quality of generated images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2022). To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth bboxes. The average of these maximum IoUs is called mIOU. In addition, Grounding-DINO AP scores are also reported.
@@ 5.3. Comparison
Compare with automatic pipelines for advertising image design. Since there are no public methods for advertising image design conditioned on the product foreground, taglines, and target size, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation. We also finetune the SDXL-based inpainting model on our PIL dataset. The only difference between this method and ours is the absence of layout control and the corresponding training stage. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general aesthetic rules. Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison. PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 5). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded tagline boxes, making them hard to read, while our model provides more flexible and readable layouts.
@@
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score. This means that our model effectively aligns with the e-commerce prompt dataset distribution after training. We also visualize the feature distribution of prompts generated in the appendix.
Table 2. Evaluation of prompt generation methods.
@@
Compare with layout generation methods. To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a). For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased/generated set. The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased areas. The generated set contains images re-generated using their ground-truth prompts (image captions). The generated set has no above hints and conforms to real applications. As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and foreground. Our method uses the ground-truth prompt, product foreground, and taglines for prediction. The results are shown in Table 3 and Figure 6. Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements. This may be due to the fixed background image content, leaving less room for the model flexibility. RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating tagline boxes that cannot adequately fill the tagline content. The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images. Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
Table 3. Quantitative performance comparison of layout generation methods. The bold number represents the best result in each column except for the erased set. BG/FG-based are the abbreviations for Background/Foreground-based methods.
@@
Figure 6. Qualitative comparison of layout generation methods. Orange: tagline, Yellow: underlay, Red: logo, Black: Invalid.
@@
5.4. Online Result
To assess the online performance of PAID, we conduct A/B tests in two advertising recommendation scenarios on Taobao. We randomly select 5,000 products from various categories (including clothing, cosmetics, food, and electronics) and collect their foreground images and taglines from the advertisers. We then generate advertising images for each product using both PAID and Pipeline 2 (the top-performing baseline as shown in Table 1). The A/B tests are conducted using 5% of the main traffic, affecting only the experimental products when this traffic is directed to them. After gathering data over one month, we found that PAID achieved a 3.02% and 3.03% increase in click-through rate (CTR) for the two recommendation scenarios, respectively. This indicates that PAID is capable of producing more visually appealing advertising images, resulting in improved recommendation outcomes.
6. Conclusion
In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create advertising images only based on product foreground images, marketing taglines, and target sizes. PAID consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates adaptive prompts according to the product foregrounds. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is then applied to get the final images. We train separate experts to conduct these sub-tasks. Two datasets are created for the convenience of training and testing. The results of the test set and the online A/B tests have proven that the proposed PAID framework generates more visually pleasing and attractive advertising images.
Acknowledgements
Appendix A Dataset Examples
In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 8 and 9, respectively. The PITA dataset includes advertising images with marketing taglines. In addition to depicting the layout of both graphic and nongraphic elements as shown in the figures, we also label the image caption (prompt), tagline content, and product foreground masks to train the prompt and layout generation models. Conversely, the PIL dataset contains product images without taglines. We annotate the layout of nongraphic elements and product foreground masks for training the image generation model.
Figure 8. Examples from the PITA dataset, with tagline, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
Figure 9. Examples in PIL dataset.
Appendix B Visualization of Prompt Distribution
We use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 10). This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
Appendix C Ablation Studies
To verify the effectiveness of our design, we conduct ablation studies on the methods and strategies used in each module.
Figure 10. T-sne visualization of generated prompts.
C.1. Effect of JPGNL in Layout Generation
Our method predicts the overall layout of graphic and nongraphic elements. Here, we examine the need to predict the layout of nongraphic elements according to the background prompt. As shown in Figure 11, without specifying the position of nongraphic elements, taglines placed on complex areas can lead to a messy appearance. By predicting the overall layout, we improve the tagline readability of advertising images. Additionally, with nongraphic layout predictions, generated images align more closely with background prompts. Without JPGNL, images might miss some objects from the prompts. Besides, in Table 5, we quantitatively analyze the impact of predicting nongraphic layouts on graphic layouts. Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and tagline readability. However, since it adds complexity to the task, it negatively affects graphic metrics such as element overlap and alignment.